OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y x - 5 y -x - 4 (-5, 2) (5, -2) (-5, -2) (5, 2)

OpenStudy (anonymous):

(5, -2)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

do you think you ca explain how you got that?

OpenStudy (anonymous):

i'm having trouble understanding it

OpenStudy (anonymous):

@adrianajones

OpenStudy (anonymous):

yes sure, okay so we have: y x - 5 and y -x - 4 let's substitute each point and check a)5,2 in y<=x-5 2<=5-5 => 2<=0. Not true meaning it's not the answer. b) 5,-2 in -2<=5-5 -2<=5-5 -2<=0. This is true. y>=-x-4 -2>=-5-4. True c) (-5,-2) y<=x-5 -2<=-5-5 -2<=-10. Not true meaning it's not the answer. d)(-5,2) 2<=-5-5 2<=-10. Not true meaning it's not the answer.

OpenStudy (anonymous):

hope you understand it now

OpenStudy (anonymous):

thanks